# To Shuffle or not to Shuffle: Auditing DP-SGD with Shuffling

### v2

@@ 9 Conclusion
This paper presented the first audit of DP-SGD implementations that use shuffling but report DP guarantees as though Poisson sub-sampling was used, investigating the differences between the theoretical guarantees of the latter and the empirical privacy leakage from the former. We did so as using shuffling to sample batches in DP-SGD has become common practice [49, 19, 40] as it is much more computationally efficient.
We introduced new auditing techniques and audited DP-SGD (Shuffle) under varying parameter settings and threat models. Our experiments revealed, even in real-world settings, a substantial gap between the empirical privacy leakage observed from DP-SGD (Shuffle) and the theoretical guarantees of DP-SGD (Poisson). More precisely, we experimented with 101 settings from SOTA models using shuffling [19, 40], finding substantial gaps in two-thirds of them. We also discussed how these gaps can depend on the batch size used and the threat model, with smaller batch sizes and stronger threat models resulting in larger gaps.
@@
Finally, we showed that our auditing techniques can be used to detect the presence of bugs in the implementations of the shuffling procedure. We investigated two common bugs [49] found in public code repositories of non-private model training, showing that our auditing framework effectively identifies them and detects even higher empirical privacy leakages in the presence of these bugs.
Implications. The measurable difference between the empirical privacy leakage observed from DP-SGD (Shuffle) and the theoretical guarantees of DP-SGD (Poisson) shows that prior work [19, 40] substantially overestimated the privacy guarantees of SOTA private models. In some instances, the actual privacy leakage is more than 4$\times$ greater than the theoretical DP guarantees.
Consequently, our work calls into question the validity of theoretical guarantees and prompts the need to formally analyze the privacy guarantees provided by shuffling the dataset. Our analysis also attests to the dangers of making seemingly simple modifications to provably correct DP algorithms.
Threat Models. Our analysis considers active white-box threat models when auditing DP-SGD (Shuffle). Specifically, we assume that the adversary can insert a target sample of their choice and view the final trained model and that they can insert (varying sizes of) canary gradients at each step and observe all intermediate models. Although the adversaries may not always have these capabilities in real-world settings, we emphasize that the provably correct DP guarantees are meant to be robust against worst-case adversaries, making them valid in the context of auditing. In fact, the same assumptions were made in prior DP-SGD audits [44, 43].
Interestingly, we detected significant differences in empirical privacy leakages depending on the threat models we experimented with. This suggests that the empirical privacy leakage for DP-SGD (Shuffle) may be lower under weaker/more realistic threat models – similar to findings from prior work for DP-SGD (Poisson) [29, 15, 59, 45]. Therefore, investigating whether the empirical privacy leakage from DP-SGD (Shuffle) exceeds the guarantees provided by DP-SGD (Poisson) even under weaker threat models would be an interesting future direction.
@@
Future Work. In future work, we plan to minimize the number of training runs required to audit DP-SGD (Shuffle) accurately. Recent work presented methods to audit DP-SGD (Poisson) with just one training run [2, 52, 48, 42]. However, these techniques may not directly apply to DP-SGD (Shuffle) and currently underestimate the privacy leakage even under power threat models, requiring substantial effort to adapt to our setting. Furthermore, our audits are currently weaker for larger batch sizes, which we believe is mainly due to the bias introduced by the “other” samples. Therefore, we also plan to reduce this bias term using debiasing techniques that can enable our audits to be effective even for large batch sizes.
@@ Appendix A Differentially Private Stochastic Gradient Descent (DP-SGD)
Appendix B Estimating $\varepsilon_{emp}$ from scores
Algorithm 5 Estimating $\varepsilon_{emp}$ from scores
1:Scores from $\mathcal{M}(D)$, $\mathcal{S}$. Scores from $\mathcal{M}(D^{\prime})$, $\mathcal{S}^{\prime}$. Significance level, $\alpha$. Privacy parameter, $\delta$. $\triangleright$ Assume $|\mathcal{S}|=|\mathcal{S}^{\prime}|$.
2:$R\leftarrow|\mathcal{S}|$triangleright Initial FPR and FNR for classifying all scores as $s\sim\mathcal{M}(D)$.
3:$\text{FPR}\leftarrow R$
4:$\text{FNR}\leftarrow 0$
5:for $\tau\in\text{sort_increasing}(\mathcal{S}\cup\mathcal{S}^{\prime})$ do $\triangleright$ Calculate FPR and FNR for classifying score as $s\sim\mathcal{M}(D)$ if $s>\tau$.
6:    if $\tau\in S$ then
7:         $\text{FNR}\leftarrow\text{FNR}+1$
8:    else
9:         $\text{FPR}\leftarrow\text{FPR}-1$     $\triangleright$ Calculate confidence intervals for FPR and FNR.
10:    $\overline{\text{FPR}}\leftarrow\text{Clopper-Pearson}(FPR,R,\alpha)$
11:    $\overline{\text{FNR}}\leftarrow\text{Clopper-Pearson}(FNR,R,\alpha)$triangleright Estimate $\varepsilon_{emp}$ from Equation 1
12:    $\varepsilon_{emp}[\tau]=\max\left\{\ln\left(\frac{1-\overline{\text{FPR}}-\delta}{\overline{\text{FNR}}}\right),\ln\left(\frac{1-\overline{\text{FNR}}-\delta}{\overline{\text{FPR}}}\right),0\right\}$
13:return $\max_{\tau}\varepsilon_{emp}[\tau]$.
As discussed in Section 3.3, we audit mechanisms by assigning a scalar “score” to each observation and then find the optimal threshold maximizing the empirical estimates $\varepsilon_{emp}$ from these scores as done in prior work [3, 11, 32, 44, 43, 48, 61]. However, we substantially improve the efficiency of this procedure by first sorting the scores in increasing order, as reported in Algorithm 5. This enables us to find the optimal threshold from, potentially, billions of observations without incurring a prohibitive computational cost.
Appendix C Model Architectures
@@
Figure 10: Comparison between a small dataset with eight records being fully shuffled vs. partially shuffled with a buffer of $K{=}4$ and batch size $B{=}2$. The target record $x_{T}$ is highlighted in red.
Appendix D Visualizing Shuffle Implementations
D.1 Partial Shuffling
In Figure 11, we visualize the possible shufflings of a small dataset with eight records – including a “target record” $x_{T}$ – when fully shuffled and when partially shuffled using a buffer $K=4$. We set batch size $B=2$, which yields four batches in each run. When fully shuffled, $x_{T}$ appears randomly in all four batches across multiple runs. However, when partially shuffled, $x_{T}$ can only appear in the first two batches. This means that, in practice, when partially shuffled, the remaining two batches do not provide any extra information to the adversary and can be discarded. Therefore, to audit partial shuffling, the adversary makes “guesses” on the size of the buffer and discards the remaining batches when running the distinguishing function.
D.2 Batch-then-Shuffle
To illustrate the Batch-then-Shuffle setting, we first consider the following one-dimensional dataset $D=(x_{T},-1,-1,-1,+1,+1,+1,+1)$ where $x_{T}$ is the target record. If this dataset is first shuffled and then batched with batch size $B=4$, each individual ‘-1’ at the beginning and each ‘+1’ at the end of the dataset can appear in either the first or second batch (nearly) independently. However, if the dataset is first batched and the batches are then shuffled, the ‘-1’s at the beginning of the dataset are always grouped together, and similarly, the ‘+1’s at the end. This makes the batch containing the target record much more easily identifiable, as visible from Figure 11.

### v3

@@ 9 Discussion and Conclusion
This paper presented the first audit of DP-SGD implementations that shuffle the training data and deterministically iterate over fixed-size batches, which we denoted as DP-SGD (Shuffle). Arguably, this addressed an important research problem as using shuffling to sample batches in DP-SGD has become common [47, 19, 39], mostly due to better computational efficiency, even though theoretical guarantees are reported as if Poisson subsampling was used, which we referred to as DP-SGD (Poisson). This makes it crucial to investigate the gap between the empirical privacy leakage from the former and the theoretical guarantees of the latter.
@@
We introduced new auditing techniques and audited DP-SGD (Shuffle) with different parameter settings and threat models. More precisely, we experimented with 101 settings from SOTA models using shuffling [19, 39], finding gaps in two-thirds of them—in some cases, the empirical privacy leakage from DP-SGD (Shuffle) was up to 4 times higher than the theoretical guarantee from DP-SGD (Poisson). We also showed how these discrepancies can depend on the batch size and the strength of the adversary, with smaller batch sizes and stronger threat models yielding larger gaps. Finally, we used our techniques to detect the presence of bugs in the implementations of the shuffling procedure. We investigated two common bugs [47] found in public code repositories of non-private model training, showing that our auditing framework effectively identifies them and detects even higher empirical privacy leakages in their presence.
Implications and Recommendations for Practitioners. The measurable differences between the empirical and the theoretical guarantees shows that prior work [19, 39] substantially overestimated the privacy guarantees of some SOTA private models. Consequently, our work calls into question the validity of reporting theoretical guarantees using DP-SGD (Poisson) while implementing algorithms that use DP-SGD (Shuffle), reiterating the importance of formally analyzing the privacy guarantees provided by shuffling the dataset. Concretely, we recommend practitioners avoid shuffling until tighter guarantees are proven. Recently proposed alternative sampling schemes like Balls-in-Bins [13, 15] could also be used, as they could balance shuffling’s computational efficiency with provably tight guarantees.
Adversarial Models. Our audits use adversaries in active white-box models; specifically, we assume they can insert a target sample of their choice and view the final trained model, and can insert (varying sizes of) canary gradients at each step and observe all intermediate models. While we introduce an adversarial model unique to shuffling, i.e., Partially-Informed, this serves as a hypothetical middle-ground between the Worst-Case adversary and a weaker one (Target-Canary), and may not necessarily reflect an actual adversary. Nevertheless, as discussed in Section 5.1, this primarily serves as a useful auditing tool, enabling us to identify the impact that the bias from other (non-target) samples has on our novel likelihood-based auditing method.
Overall, although not all the adversaries we consider may always have the corresponding capabilities against models deployed in production, recall that provably correct DP guarantees are meant to be robust in worst-case settings, making our adversaries valid in the context of auditing. In fact, similar assumptions were made in prior DP-SGD audits [43, 42, 4].
@@
Limitations & Future Work. One main limitation of our work is the computational cost required to run the audit. Specifically, since shuffling does not provide any improved $f$-DP or $\mu$-GDP guarantees, we require training runs in the order of thousands and millions. Recent work presented audits of DP-SGD (Poisson) with just one run [2, 50, 46, 41]; however, these techniques may not directly apply to DP-SGD (Shuffle) and underestimate privacy leakage even under powerful threat models. Thus, it is not yet clear how to adapt them to our setting. In follow-up work, we plan to investigate how to minimize the number of training runs required to audit DP-SGD (Shuffle) accurately, which would also enable us to audit larger deep learning models.
In addition, our audits are weaker for larger batch sizes, which we believe is mainly due to the bias introduced by the “other” samples. We plan to reduce this bias term using debiasing techniques (e.g., [51]) to make our audits effective even for large batch sizes.
Finally, due to computational constraints, we only audited one epoch of DP-SGD (Shuffle) and therefore did not measure model utility. In the future, we plan to audit state-of-the-art models built using DP-SGD (Shuffle) directly, which will also allow us to measure drops in utility with different adversaries. As mentioned, we expect utility to plummet with Worst-Case but remain stable with Partially-Informed and Target-Canary adversaries.
Acknowledgements. This work has been supported by the National Science Scholarship (PhD) from the Agency for Science Technology and Research, Singapore.
@@ Appendix A Differentially Private Stochastic Gradient Descent (DP-SGD)
Appendix B Worst-Case vs Pathological Dataset
We now briefly explain how our Worst-Case adversary can be modeled with the Pathological Dataset adversary used in prior work [43]. The Pathological Dataset adversary crafts the malicious dataset by making sure the dataset (not including the target sample) is labeled perfectly by the initial model. This ensures that the gradients of all other samples are zero, except for the target sample. Furthermore, to ensure that subsequent model updates do not disrupt the gradients, the Pathological Dataset also sets the learning rate to 0.
On the other hand, in our work, the Worst-Case adversary crafts malicious gradients for all samples that are equal in magnitude but opposite in direction to the target sample. This can be modeled by the Pathological Dataset adversary by first calculating the gradient of a given target sample with respect to the initial model parameters. Then, the other samples can be crafted by optimizing an input sample with respect to the target sample’s gradient in the opposite direction and repeating them $N$ times. Finally, the learning rate can be set to 0 to ensure that future model updates do not disrupt the gradients. Therefore, we can ensure that the gradient of the target sample is always equal in magnitude and opposite in direction to the target sample throughout the training process.
Appendix C Model Architectures